The punk package^{∗} for use with L^{A}TEX2e

Gerd Neugebauer Mainzer Str. 16 56321 Rhens

             gerd@informatik.uni-koblenz.de

1996/05/19

1 Introduction

The font punk provides a punkphic font. This package provides means to use this font.

2 Usage

This file can be used as a package by placing its name in the argument of \usepackage. Afterwards the font family punk is defined. This could also have been done by providing a font definition file.

\punk The command \punk changes the current font family to punk and the encoding to T1. Usually this should be used in a TEX group only since it changes the font encoding.

3 Implementation

First we declare a new font family for the punk font.

1 \DeclareFontFamily{OT1}{punk}{}

This font is only available in the normal shape. Here we can get the desired font by (silently) scaling the only present fonts.

2 \DeclareFontShape{OT1}{punk}{m}{n}{<->s*punk20}{}
3 \DeclareFontShape{OT1}{punk}{m}{sl}{<->s* punksl20}{}
4 \DeclareFontShape{OT1}{punk}{m}{it}{<->sub*punk/m/sl}{}
5 \DeclareFontShape{OT1}{punk}{bx}{n}{<->s*punkbx20}{}
6 \DeclareFontShape{OT1}{punk}{b}{n}{<->sub*punk/bx/m}{}
7 \DeclareFontShape{OT1}{punk}{bx}{sl}{<->sub*punk/bx/m}{}
8 \DeclareFontShape{OT1}{punk}{bx}{it}{<->sub*punk/bx/m}{}
9 \DeclareFontShape{OT1}{punk}{b}{sl}{<->sub*punk/bx/m}{}
10 \DeclareFontShape{OT1}{punk}{b}{it}{<->sub*punk/bx/m}{}

∗This file has version number v1.1, dated 1996/05/19.

Now we define font changing command.

\punk The macro \punk selects the punk family.

11 \newcommand\punk{%
12 \fontfamily{punk}%
13 \fontencoding{OT1}%
14 \selectfont}

\textpunk The macro \textpunk typesets its argument in the punk font.

15 \newcommand\textpunk[1]{\begingroup
16 \fontfamily{punk}\fontencoding{OT1}\selectfont
17 #1\endgroup}